OpenStudy (ammarah):

Write a polynomial function in standard form with zeros at 2,-1, and 3.

OpenStudy (tkhunny):

A zero of 'a' corresponds to a factor of (x-a). Go!

OpenStudy (ammarah):

? i dont get it

ganeshie8 (ganeshie8):

since 2 is a zero, (x-2) is a factor of polynomial

ganeshie8 (ganeshie8):

so you can write polynomial function as : f(x) = (x-2)(x--1)(x-3)

ganeshie8 (ganeshie8):

multiply and put it in standard form : ax^3 + bx^2 + cx + d

OpenStudy (ammarah):

i got x^3-6x^2+10x-6

OpenStudy (ammarah):

@ganeshie8 is this correct^ ???????

ganeshie8 (ganeshie8):

nope. check wid http://www.wolframalpha.com/input/?i=+%28x-2%29%28x--1%29%28x-3%29

OpenStudy (ammarah):

i dont get how to do it tho

OpenStudy (tkhunny):

You were just shown the entire setup. Please look carefully at what is already written. Demonstrate your best efforts. So far, you have told us only that you don't get it. Please read carefully. Show your work. It's all right here.

OpenStudy (ammarah):

oh...

OpenStudy (ammarah):

i kinda dont get it cuz all im doing is multiplying

OpenStudy (tkhunny):

Pretty much it after you find the factors. Good work.

OpenStudy (ammarah):

? i got x^3-6x^2+7x+6

OpenStudy (tkhunny):

Is that right?

OpenStudy (ammarah):

idk im asking u

OpenStudy (tkhunny):

If I say "yes" or "no", will it help at all? You need to gain the confidence to tell me it is right. Is it. There were three linear factors. Did it come out x^3? That would be a good sign. The constant terms are -2, +1, -3. Did the final constant term come out +6? That would be a good sign.

OpenStudy (ammarah):

ok

OpenStudy (tkhunny):

In this case, despite those two easy checks, the other two are no good. Please try again. Do the first two binomials and show me what you get.

OpenStudy (ammarah):

x^2-3x-2

OpenStudy (tkhunny):

There's your problem. -3 is not correct. Please try again.

OpenStudy (ammarah):

x^2-x-2

OpenStudy (tkhunny):

That's better. Now introduce the 3rd factor. (x^2 - x - 2)(x-3) = ??

OpenStudy (ammarah):

x^3-4x^2-x+6

OpenStudy (tkhunny):

So close! -2x + 3x = +1x

OpenStudy (ammarah):

oh oops thanks can u help me with some more please questions

OpenStudy (tkhunny):

If you're done with this one, please close it and start a new thread.